# Eccentric binary black holes: A new framework for numerical relativity waveform surrogates

### v1

@@ Introduction:
The relaxation of the quasi-circular restriction for binary systems enlarges the parameter space by two dimensions, and introduces a new characteristic frequency to the gravitational wave signal (corresponding to the radial motion of the binary). Recently, there has been a significant effort in the development of waveform models incorporating this additional physics Ramos-Buades et al. (2023a); Gamboa et al. (2024a); Liu et al. (2023); Nagar et al. (2024); Gamba et al. (2024); Paul et al. (2025); Planas et al. (2025a); Albanesi et al. (2025); Gamboa et al. (2024b); Morras (2025); Henry and Khalil (2023); Konigsdorffer and Gopakumar (2006); Morras et al. (2025a); Setyawati and Ohme (2021); Chattaraj et al. (2022); Manna et al. (2025); Islam (2024); Islam and Venumadhav (2025a); Islam et al. (2025a). With these models in hand, several works have begun to perform searches for eccentric signals in data provided by LIGO, Virgo, and KAGRA Wang and Nitz (2021); Nitz et al. (2019); Abbott et al. (2019b); Abac et al. (2024b); Phukon et al. (2025), as well as re-analyzing previously detected events to search for eccentricity Morras et al. (2025b); Gupte et al. (2024); Clarke et al. (2022); Ramos-Buades et al. (2023a); Planas et al. (2025b); Kacanja et al. (2025). However, there are caveats; the increased computational cost of these models, as well as the additional dimensionality introduced by eccentricity, currently prohibits the usage of these models in routine parameter estimation pipelines Kacanja et al. (2025); Gupte et al. (2024). Along with this, these models do not include complete eccentric corrections to the merger and ringdown phases, and so their ability to capture eccentric signatures for very high-mass systems is limited.
A natural candidate to alleviate these issues is the surrogate family of waveform models Field et al. (2014); Varma et al. (2019a, b); Blackman et al. (2017, 2015); Islam et al. (2021a, 2022); Gadre et al. (2022). Surrogates rely on a data-driven, reduced order modelling approach, where one interpolates between a set of basis waveforms. A typical choice of basis waveforms are those resulting from numerical relativity (NR) simulations Baumgarte and Shapiro (2010); Pretorius (2005); Campanelli et al. (2006); Baker et al. (2006); Mroue et al. (2013); Boyle et al. (2019); Scheel et al. (2025). NR involves fully simulating Einstein’s equations with no underlying approximations Baumgarte and Shapiro (2010), and so captures all of the relevant physics of the system. These waveforms, while costly, serve as the “ground truth” for evaluating the precision of approximate waveform models, and are routinely used for calibration and improvement of said models (e.g. Pompili et al. (2023, 2025); Pratten et al. (2020)). Surrogate models, when trained directly on NR waveforms, are able to faithfully reproduce the underlying waveforms to within their intrinsic errors, providing us with a model that is both fast and accurate.
@@
In order to re-parameterize the system as a function of radial phase, we need an approach to extract a radial phase from numerical relativity simulations. Several definitions of radial phase could be chosen: for this work, we employ the relativistic anomaly Gamboa et al. (2024b), which we obtain by solving the following quasi-Keplerian equations:
@@
where $e$ corresponds to the quasi-Keplerian eccentricity, $\zeta$ the relativistic anomaly, $\nu=m_{1}m_{2}/M^{2}$ the symmetric mass-ratio, $M$ the total mass of the binary, $\Omega$ is the instantaneous orbital frequency of the binary, and $x$ is the dimensionless orbit-averaged orbital frequency. The full 33PN equations can be found in Gamboa et al. (2024a). This radial phase is desirable, as (unlike the mean anomaly used in previous work Islam et al. (2021b)) it is both a smooth function of time, and has a well-defined limit for $e\rightarrow 0$. These equations are currently derived up to 3PN for aligned spin systems, and up to 2PN for generic systems.
To solve these equations, one must supply $(q,\chi_{i},\Omega(t),e_{0},\zeta_{0})$. Since we are restricted to non-spinning systems, we set $\chi_{i}=0$, and compute the mass-ratio $q$ using the Christodolou mass computed from the apparent horizons Scheel et al. (2025). For $\Omega(t)$, we use the orbital frequency computed using the coordinate center’s of the individual black-holes.
@@
Our goal is to generate an “inspiral” surrogate that is applicable for times up to t=1200Mt=-1200M. To this end, we identify the first periastron passage that occurs after t=1200Mt=-1200M for each waveform in the dataset, and restrict the $\zeta$-parameterisation up to this point. Due to length limitations of the underlying NR waveforms, we consider the $11$ radial periods immediately before this for modelling. In addition to modelling $A_{22}(\zeta)$ and $\phi_{22}(\zeta)$, we will also model $t(\zeta)$, so that we can transform the surrogate evaluation back to the time domain. Finally, for only the inspiral surrogate, we perform a rotation and time-shift to each waveform, such that $t=0$ and $\phi_{22}=0$ at the final periastron passage to be modelled, to further simplify the data.
@@ Results:
In order to demonstrate the efficacy of this decomposition, we construct a surrogate for the $(2,2)$ mode for non-spinning eccentric systems. We use a training set of $156$ waveforms from the SXS catalog Scheel et al. (2025); Nee et al. (2025), with identifiers SXS:BBH:[3731-3747, 3749-3776, 3778-3820, 4293-4304, 4321-4332, 4357-4381, 4442-4455]. The training set consists of mass-ratios $q\in[1,4]$ and eccentricities $e_{-3000M}\in[10^{-3},0.25]$. For fits across parameter space, we employ Gaussian process regression as in Varma et al. (2019c); Taylor and Varma (2020). The model, which we refer to as NRSurE_q4NoSpin_22, will be made publicly available as part of the GWSurrogate package Field et al. (2025) in the near future.
@@
To validate the model, we perform a leave-eight-out cross validation study, where we divide the training set randomly into groups of 8 cases each. We build a series of surrogates, each excluding one group of waveforms from their training set, which is then validated against the excluded waveforms. The effectiveness of this modelling approach is demonstrated in the right panel of Fig. 4. Plotted are the mismatches from the leave-eight-out study, along with the mismatches obtained comparing the two highest resolutions for each system, which serves as a conservative estimate for the accuracy of the training waveforms. We see that the surrogate is able to faithfully reproduce the underlying numerical relativity waveforms, reducing the largest mismatches by several orders of magnitude, and the median mismatch by a factor of $\sim 4$. This indicates that the limiting factor in the accuracy of the surrogate is likely the accuracy of the underlying numerical relativity waveforms, and not the surrogate modelling techniques (similar to what was seen in Varma et al. (2019a)). As an extra illustration of the accuracy of the model, the left panel of Fig. 4 shows the worst validation case waveform, corresponding to one of the largest mass-ratios and eccentricities in the training set (i.e. a boundary case). The inclusion of systems with larger eccentricities and mass-ratios would likely improve the accuracy of the model for this case.
@@ Conclusion:
While the $(2,2)$ mode predicted by the model is accurate to within the errors of the underlying NR waveforms, it is still restricted in its length. This can be alleviated by hybridization of the NR waveforms with analytical approximants Sun et al. (2024); Varma et al. (2019b). In particular, waveform models that already solve the quasi-Keplerian equations (such as SEOBNRv5EHM Gamboa et al. (2024a)) would be a natural choice. Regardless, by ensuring that future simulation campaigns focus on length in terms of radial periods as opposed to frequency/time, we can ensure that the surrogate is sufficiently long for most applications.
@@ Acknowledgments:
The authors would like to thank Arnab Dhani, Aldo Gamboa, Marcus Haberland, Benjamin Leather, Oliver Long, Philip Lynch, and Lorenzo Pompili for useful discussions, as well as Keefe Mitman and Lucy Thomas for feedback on a draft of this work. P.J.N. would like to thank Aurora Abbondanza for one-way conversations regarding this work. T.I. is supported in part by the National Science Foundation under Grant No. NSF PHY-2309135 and the Simons Foundation (216179, LB). A. R.-B. is supported by the Veni research programme which is (partly) financed by the Dutch Research Council (NWO) under the grant VI.Veni.222.396; acknowledges support from the Spanish Agencia Estatal de Investigación grant PID2022-138626NB-I00 funded by MICIU/AEI/10.13039/501100011033 and the ERDF/EU; is supported by the Spanish Ministerio de Ciencia, Innovación y Universidades (Beatriz Galindo, BG23/00056) and co-financed by UIB. This material is based upon work supported by the National Science Foundation under Grants No. PHY-2309211, No. PHY-2309231, and No. OAC-2209656 at Caltech; by No. PHY-2407742, No. PHY-2207342, and No. OAC-2209655 at Cornell; and by No. PHY-2208014 and AST-2219109 at Cal State Fullerton. This work was supported by the Sherman Fairchild Foundation at Caltech and Cornell, and by the Dan Black Family Trust, and Nicholas and Lee Begovich at Cal State Fullerton. S.F. acknowledges support from NSF Grants No. AST-2407454 and PHY-2110496. H. R. R. acknowledges financial support provided under the European Union’s H2020 ERC Advanced Grant “Black holes: gravitational engines of discovery” grant agreement no. Gravitas–101052587. Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. V.V. acknowledges support from NSF Grant No. PHY-2309301. This work was partly supported by UMass Dartmouth’s Marine and Undersea Technology (MUST) research program funded by the Office of Naval Research (ONR) under grant No. N00014-23-1-2141. The computations presented here were conducted in the Resnick High Performance Computing Center, a facility supported by Resnick Sustainability Institute at the California Institute of Technology. This work used Expanse at San Diego Supercomputer Center Strande et al. (2021), Stampede 2 at Texas Advanced Computing Center, through allocation PHY990002 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by U.S. National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296 Boerner et al. (2023). The authors acknowledge the Texas Advanced Computing Center (TACC) at The University of Texas at Austin for providing computational resources that have contributed to the research results reported within this paper. URL: http://www.tacc.utexas.edu. This work used the Extreme Science and Engineering Discovery Environment (XSEDE), which is supported by National Science Foundation grant number ACI-1548562. Specifically, it used the Bridges-2 system, which is supported by NSF award number ACI-1928147, at the Pittsburgh Supercomputing Center (PSC). Computations were performed on the Urania cluster at the Max Planck Computing and Data Facility.

### v2

@@ Introduction:
The relaxation of the quasi-circular restriction for binary systems enlarges the parameter space by two dimensions, and introduces a new characteristic frequency to the gravitational wave signal (corresponding to the radial motion of the binary). Recently, there has been a significant effort in the development of waveform models incorporating this additional physics Ramos-Buades et al. (2023a); Gamboa et al. (2025); Liu et al. (2024); Nagar et al. (2024); Gamba et al. (2024); Paul et al. (2025); Planas et al. (2025a); Albanesi et al. (2025); Gamboa et al. (2024); Morras (2025); Henry and Khalil (2023); Konigsdorffer and Gopakumar (2006); Morras et al. (2025a); Setyawati and Ohme (2021); Chattaraj et al. (2022); Manna et al. (2025); Islam (2024); Islam and Venumadhav (2025a); Islam et al. (2025a). With these models in hand, several works have begun to perform searches for eccentric signals in data provided by LIGO, Virgo, and KAGRA Wang and Nitz (2021); Nitz et al. (2019); Abbott et al. (2019b); Abac et al. (2024b); Phukon et al. (2025), as well as re-analyzing previously detected events to search for eccentricity Morras et al. (2025b); Gupte et al. (2024); Clarke et al. (2022); Ramos-Buades et al. (2023a); Planas et al. (2025b); Kacanja et al. (2025); Jan et al. (2025). However, there are caveats; the increased computational cost of these models, as well as the additional dimensionality introduced by eccentricity, currently prohibits the usage of these models in routine parameter estimation pipelines Kacanja et al. (2025); Gupte et al. (2024). Along with this, these models do not include complete eccentric corrections to the merger and ringdown phases, and so their ability to capture eccentric signatures for very high-mass systems is limited.
A natural candidate to alleviate these issues is the surrogate family of waveform models Field et al. (2014); Varma et al. (2019a, b); Blackman et al. (2017, 2015); Islam et al. (2021a, 2022); Gadre et al. (2022). Surrogates rely on a data-driven, reduced order modelling approach, where one interpolates between a set of basis waveforms. A typical choice of basis waveforms are those resulting from numerical relativity (NR) simulations Baumgarte and Shapiro (2010); Pretorius (2005); Campanelli et al. (2006); Baker et al. (2006); Mroue et al. (2013); Boyle et al. (2019); Scheel et al. . NR involves fully simulating Einstein’s equations with no underlying approximations Baumgarte and Shapiro (2010), and so captures all of the relevant physics of the system. These waveforms, while costly, serve as the “ground truth” for evaluating the precision of approximate waveform models, and are routinely used for calibration and improvement of said models (e.g. Pompili et al. (2023, 2025); Pratten et al. (2020)). Surrogate models, when trained directly on NR waveforms, are able to faithfully reproduce the underlying waveforms to within their intrinsic errors, providing us with a model that is both fast and accurate.
@@
In order to re-parameterize the system as a function of radial phase, we need an approach to extract a radial phase from numerical relativity simulations. Several definitions of radial phase could be chosen: for this work, we employ the relativistic anomaly Gamboa et al. (2024), which we obtain by solving the following quasi-Keplerian equations:
@@
where $e$ corresponds to the quasi-Keplerian eccentricity, $\zeta$ the relativistic anomaly, $\nu=m_{1}m_{2}/M^{2}$ the symmetric mass-ratio, $M$ the total mass of the binary, $\Omega$ is the instantaneous orbital frequency of the binary, and $x$ is the dimensionless orbit-averaged orbital frequency. The full 33PN equations can be found in Gamboa et al. (2025). This radial phase is desirable, as (unlike the mean anomaly used in previous work Islam et al. (2021b)) it is both a smooth function of time, and has a well-defined limit for $e\rightarrow 0$. These equations are currently derived up to 3PN for aligned spin systems, and up to 2PN for generic systems.
To solve these equations, one must supply $(q,\chi_{i},\Omega(t),e_{0},\zeta_{0})$. Since we are restricted to non-spinning systems, we set $\chi_{i}=0$, and compute the mass-ratio $q$ using the Christodolou mass computed from the apparent horizons Scheel et al. . For $\Omega(t)$, we use the orbital frequency computed using the coordinate center’s of the individual black-holes.
@@
Our goal is to generate an “inspiral” surrogate that is applicable for times up to t=1200Mt=-1200M. To this end, we identify the first periastron passage that occurs after t=1200Mt=-1200M for each waveform in the dataset, and restrict the $\zeta$-parameterisation up to this point. Due to length limitations of the underlying NR waveforms, we consider the $11$ radial periods immediately before this for modelling. In addition to modelling $A_{22}(\zeta)$ and $\phi_{22}(\zeta)$, we also model $t(\zeta)$, so that we can transform the surrogate evaluation back to the time domain. Finally, for only the inspiral surrogate, we perform a rotation and time-shift to each waveform, such that $t=0$ and $\phi_{22}=0$ at the final periastron passage, to further simplify the data.
@@ Results:
In order to demonstrate the efficacy of this decomposition, we construct a surrogate for the $(2,2)$ mode for non-spinning eccentric systems. We use a training set of $156$ waveforms from the SXS catalog Scheel et al. ; Nee et al. (2025), with identifiers SXS:BBH:[3731-3747, 3749-3776, 3778-3820, 4293-4304, 4321-4332, 4357-4381, 4442-4455]. The training set consists of mass-ratios $q\in[1,4]$ and eccentricities $e_{-3000M}\in[10^{-3},0.25]$. For fits across parameter space, we employ Gaussian process regression as in Varma et al. (2019c); Taylor and Varma (2020). The model, which we refer to as NRSurE_q4NoSpin_22, will be made publicly available as part of the GWSurrogate package Field et al. (2025) in the near future.
@@
To validate the model, we perform a leave-eight-out cross validation study, where we divide the training set randomly into groups of 8 cases. We build a series of surrogates, each excluding one group of waveforms from their training set, which is then validated against the excluded waveforms. The effectiveness of this modelling approach is demonstrated in the right panel of Fig. 4. Plotted are the mismatches from the leave-eight-out study, along with the mismatches obtained comparing the two highest resolutions for each system, which serves as a conservative estimate for the accuracy of the training waveforms. We see that the surrogate is able to faithfully reproduce the underlying numerical relativity waveforms, reducing the largest mismatches by several orders of magnitude, and the median mismatch by a factor of $\sim 4$. This indicates that the limiting factor in the accuracy of the surrogate is likely the accuracy of the underlying numerical relativity waveforms, and not the surrogate modelling techniques (similar to what was seen in Varma et al. (2019a)). As an extra illustration of the accuracy of the model, the left panel of Fig. 4 shows the worst validation case waveform, corresponding to one of the largest mass-ratios and eccentricities in the training set (i.e. a boundary case). The inclusion of systems with larger eccentricities and mass-ratios would likely improve the accuracy of the model for this case.
@@ Conclusion:
While the $(2,2)$ mode predicted by the model is accurate to within the errors of the underlying NR waveforms, it is still restricted in its length. This can be alleviated by hybridization of the NR waveforms with analytical approximants Sun et al. (2024); Varma et al. (2019b). In particular, waveform models that already solve the quasi-Keplerian equations (such as SEOBNRv5EHM Gamboa et al. (2025)) would be a natural choice. Regardless, by ensuring that future simulation campaigns focus on length in terms of radial periods as opposed to frequency/time, we can ensure that the surrogate is sufficiently long for most applications.
@@ Acknowledgments:
The authors would like to thank Arnab Dhani, Aldo Gamboa, Marcus Haberland, Benjamin Leather, Oliver Long, Philip Lynch, and Lorenzo Pompili for useful discussions, as well as Guillermo Lara, Keefe Mitman, and Lucy Thomas for feedback on a draft of this work. P.J.N. would like to thank Aurora Abbondanza for one-way conversations regarding this work. T.I. is supported in part by the National Science Foundation under Grant No. NSF PHY-2309135 and the Simons Foundation (216179, LB). A. R.-B. is supported by the Veni research programme which is (partly) financed by the Dutch Research Council (NWO) under the grant VI.Veni.222.396; acknowledges support from the Spanish Agencia Estatal de Investigación grant PID2022-138626NB-I00 funded by MICIU/AEI/10.13039/501100011033 and the ERDF/EU; is supported by the Spanish Ministerio de Ciencia, Innovación y Universidades (Beatriz Galindo, BG23/00056) and co-financed by UIB. This material is based upon work supported by the National Science Foundation under Grants No. PHY-2309211, No. PHY-2309231, and No. OAC-2209656 at Caltech; by No. PHY-2407742, No. PHY-2207342, and No. OAC-2209655 at Cornell; and by No. PHY-2208014 and AST-2219109 at Cal State Fullerton. This work was supported by the Sherman Fairchild Foundation at Caltech and Cornell, and by the Dan Black Family Trust, and Nicholas and Lee Begovich at Cal State Fullerton. S.F. acknowledges support from NSF Grants No. AST-2407454 and PHY-2110496. H. R. R. acknowledges financial support provided under the European Union’s H2020 ERC Advanced Grant “Black holes: gravitational engines of discovery” grant agreement no. Gravitas–101052587. Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. V.V. acknowledges support from NSF Grant No. PHY-2309301. This work was partly supported by UMass Dartmouth’s Marine and Undersea Technology (MUST) research program funded by the Office of Naval Research (ONR) under grant No. N00014-23-1-2141. The computations presented here were conducted in the Resnick High Performance Computing Center, a facility supported by Resnick Sustainability Institute at the California Institute of Technology. This work used Expanse at San Diego Supercomputer Center Strande et al. (2021), Stampede 2 at Texas Advanced Computing Center, through allocation PHY990002 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by U.S. National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296 Boerner et al. (2023). The authors acknowledge the Texas Advanced Computing Center (TACC) at The University of Texas at Austin for providing computational resources that have contributed to the research results reported within this paper. URL: http://www.tacc.utexas.edu. This work used the Extreme Science and Engineering Discovery Environment (XSEDE), which is supported by National Science Foundation grant number ACI-1548562. Specifically, it used the Bridges-2 system, which is supported by NSF award number ACI-1928147, at the Pittsburgh Supercomputing Center (PSC). Computations were performed on the Urania cluster at the Max Planck Computing and Data Facility.